INAF queue observing, 2023 November 18 - November 19

Observer: Felice Cusano, Simona Paiano
Support: Alex Becker
Telescope operator : David Gonzalez Huerta
Instrument: PEPSI-PFU BINO

Summary

These observations have been performed by Alex Becker with the support of LBT-Italia-obs remotely.

At sunset the sky is covered by clouds and humidity is 99%. TMS is not available. LUCI1 is not available for all the run due to technical problem.
At 01:30 UT the humidity decreased and we were able to open. Technical problem with the right shutter door while opening. Problem solved around 10:00 UT. We observe with PEPSI.

Night duration (12 degree - twilight) 01:30 - 12:40 UT = * 11h* * 10m*
Observing time: 2 h 30 min (23%)
Weather time loss: 7 h 00 min (62%)
Technical time loss: 1 h 40 min (15%)
Shark-VIS time loss: h min (%)

Data Summary

Project_RUN Objects Instrument Configuration ExpTime /Time_tot Notes
Pino 55 Cnc PEPSI CD1-CD4     S/N~300 on CD1, ~1000 on CD4
Bellazzini Gaia DR3 642798553709924992 PEPSI CD3-CD6     S/N~88 on CD3, ~95 on CD6
Bellazzini Gaia DR3 625983894184971392 PEPSI CD3-CD6     S/N~80 on CD3, ~96 on CD6
Bellazzini Gaia DR3 754423443768017280 PEPSI CD3-CD6     S/N~88 on CD3, ~102 on CD6
           
Further details in the Log by Alex Becker at this link: https://drive.google.com/file/d/1IbVbYhv9BjTpbUDZu1WoLc0flaJu6uDo/view?usp=sharing

PEPSI log: https://drive.google.com/file/d/1qu6LSIeXVF6Rf9Z583rryb3AqEmrfpyI/view?usp=sharing
Topic revision: r6 - 23 Nov 2023, FeliceCusano
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback